Solve for x, where M is molar and s is seconds. x = (7.5×10^3M^−2s^−1)(0.15M)^3
Answer #1

x = (7.5 × 103 M-2 s-1)(0.15 M)3

= 7.5 × (0.15)3 × 103 M3-2 s-1

= 0.0253125 × 103 M s-1

= 25.3125 M s-1